OpenStudy (anonymous):

[10.07] What is the exact value of the square root of 24 over 2 the square root of 64.? Simplify if possible. the square root of 6 over 8. the square root of 6 over 16. the square root of 6 over 2. the square root of 6 over 32.

OpenStudy (shamil98):

\[\huge \frac{ \sqrt{24} }{ 2\sqrt{64} }\] this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so find factors of 24 that are a perfect square so you can write it as:\[\frac{ \sqrt{4*6} }{ 2\sqrt{64} }\] then you can write that as:\[\frac{ \sqrt{4}*\sqrt{6} }{ 2\sqrt{64} }\] now simplify (the square root of 64 is 8):\[\frac{ 2\sqrt{6} }{ 16 }\] divide the top and bottom by 2: \[\frac{ \sqrt{6} }{ 8 }\]